I am a little behind please provide some explanation, will provide thumbs up.

This question considers long-run policies in Turkey relative to its largest trading partner:
Europe. Assume Turkey’s money growth rate is currently 15% and Turkey’s output growth is
9%. Europe’s money growth rate is 4% and its output growth is 3%. For the following questions,
use the conditions associated with the simple monetary model. Treat Turkey as the home
country and define the exchange rate as Turkish lira per euro, EL/ €.

a) Calculate the inflation rate in Turkey

b) Calculate the inflation rate in Europe.

c) Calculate the expected rate of depreciation in the Turkish lira relative to the euro.

d) Suppose the central bank of the Republic of Turkey decreases the money growth rate
from 15% to 11%. If nothing in Europe changes, what is the new inflation rate in Turkey?

e) Illustrate how the change in

(d) affects the following variables: MT, PT, real money
balances, and EL/ € over time.

f) Suppose the central bank of the Republic of Turkey sought to implement policy that
would cause the Turkish lira to appreciate relative to the euro. What ranges of the
money growth rate (assuming positive values) would allow the central bank to achieve
this objective? What range of values does this imply for Turkey’s inflation rate?
Answer #1

Ans

Can answer only 4 parts according to Chegg policy

1 ≠money growth-output growth=15-9=6%

2 =4-3=1%

3 6/1=6%

4 11-9=2%